# Ensemble-Based Event Camera Place Recognition Under Varying Illumination

### v1

@@ IV-B Implementation Details
We adopt Recall@1 as our primary evaluation metric, following standard practice in the VPR literature. This metric measures the ratio of correct top-1 matches to the total number of query frames:
$\text{Recall@1}=\frac{\text{Number of correct top-1 matches}}{\text{Total number of queries}}.$ (10)
To evaluate event binning strategies, we experiment with both count-based and time-based binning. All subsequent experiments use time-based binning, which consistently yields stronger performance. For each reconstructed frame, we linearly interpolate available GPS data to determine its ground-truth position. We extract features using the open-source VPR-Methods-Evaluation repository, applying a 25-metre tolerance to define correct matches [37]. This tolerance is smaller than previous methods [11, 23] for the Brisbane Event dataset, which used 70m, so our reported recall results are more conservative. No feature extractor or reconstruction model is retrained; all models are used as released.
@@ V Results and Discussion
This section presents and compares the performance of our event-based VPR system. Section I contains our evaluation of each ensemble strategy across reconstruction methods and feature extractors. Section V-B investigates the effect of our modified sequence matching approach (Section III-D) compared to the original SeqSLAM methodology. The influence of event-binning strategies on recognition accuracy and robustness is then examined in Section V-C. Finally, Section V-D compares key design choices in event-based VPR, including binning strategies, polarity, reconstruction methods, and feature extractors, critically evaluating the impact of each component on performance.
@@ V-A Event VPR via Ensembling
The effectiveness of the proposed ensemble strategies across reconstruction methods, feature extractors and the combination of feature extractors, reconstruction and temporal ensembles was evaluated against two baselines: (i) the best-performing individual ensemble member and (ii) the temporal resolution only ensemble from [11]. The results exhibit a bimodal distribution, reflecting the expected performance gap between day-day and day-night pairs due to significant domain shifts. To highlight these trends, results are grouped into day and night conditions. As shown in Figure 3, across 10 diverse traverse combinations, the proposed ensembles consistently outperformed both baselines.
Table I quantifies the relative improvement over temporal ensembling. Notably, the most significant gains occur under severe illumination variations (i.e., day to night), where the best single ensemble feature extractor achieves a 4.4% absolute and 14.5% relative improvement in Recall@1, and the combined ensembles achieve a 17.11% absolute and 56.86% relative improvement in Recall@1. Gains in daytime conditions are smaller, but still consistent. Figure 4 illustrates that the ensemble similarity matrix aligns more closely with ground-truth matches than any individual reconstruction method, yielding a 9% higher recall than the best individual in this example. This highlights the robustness gains achieved through aggregation.
TABLE I: Performance improvement over the baseline from [11].
Condition Method R@1 (%) Gain (%)
Day–Day Baseline [11] 77.16
Reconstruction 78.00 1.09
Feature Extractor 78.19 1.33
Combined Ensemble 80.13 3.85
Day–Night Baseline [11] 30.20
Reconstruction 32.79 8.58
Feature Extractor 34.58 14.51
Combined Ensemble 47.31 56.86
Figure 3: Recall@1 for each reference–query pair across ensembling strategies and best individual method, shown separately for day and night conditions. Results are averaged over sequence lengths of 10, 20, and 30 at 1 Hz sampling. The combined ensemble aggregates predictions from varied feature extractors, event-to-frame reconstructions and temporal resolutions.
@@
TABLE II: Mean Recall@1 across sequence lengths and p-values from paired t-test
@@ V-C Event Binning Strategies
Figure 5: Average Recall@1 versus the number of reference and query frames, evaluated across two binning strategies and varying event slice resolutions. Time-based binning consistently yields the highest recall, with performance improving at higher resolutions (i.e., smaller bin sizes). The total number of frames serves as a proxy for computational cost, as both event-to-frame reconstruction and feature extraction scale linearly with frame count. Consequently, higher-resolution bins incur greater compute cost for VPR.
@@ V-D Reconstructions and Feature Extractors for Event VPR
Two variants of event count reconstruction are also evaluated: with and without polarity. Including polarity improves daytime performance by 6.8% Recall@1 (52.2% vs. 45.4%), while the difference at night is minimal (11.8% vs. 11.0%). Figure 8 shows VPR performance across different feature extractors. The observed trends are consistent with those reported in standard camera-based VPR benchmarks. Earlier methods such as CosPlace and NetVLAD achieve mean Recall@1 scores of 24–25%, while more recent approaches like MixVPR and MegaLoc reach 38–42%. MegaLoc, the current state-of-the-art, performs particularly well at night, achieving 16.9% Recall@1—at least 5% higher than all other evaluated methods.
@@ VI Conclusions and Future Work
The feature extraction for the combined ensemble requires  640 ms per query, compared to  10 ms for an individual method, highlighting the trade-off between increased robustness and higher computational cost. While our study includes a diverse set of feature extractors, reconstructions, and temporal strategies, further gains may be achieved by learning event-specific feature extractors and developing reconstructions tailored for VPR in night conditions. Notably, none of the deep learning-based feature extractors used were trained on event data, suggesting a promising direction for developing event-native architectures. Future work could also explore reconstruction models that remain robust in low-light conditions, expanding the utility of event cameras in challenging environments.

### v2

@@ IV-B Implementation Details
We adopt Recall@1 as our primary evaluation metric, following standard practice in the VPR literature. This metric measures the ratio of correct top-1 matches to the total number of query frames.
To evaluate event binning strategies, we experiment with both count-based and time-based binning. All subsequent experiments use time-based binning, which consistently yields stronger performance. For each reconstructed frame, we linearly interpolate available GPS data to determine its ground-truth position. We extract features using the open-source VPR-Methods-Evaluation repository, applying a 25-metre tolerance to define correct matches [42]. This tolerance is smaller than previous methods [11, 23] for the Brisbane Event dataset, which used 70m, so our reported recall results are more conservative. No feature extractor or reconstruction model is retrained; all models are used as released. For comparison to LENS [8], we use 1 second temporal windows with event count reconstruction, a ground truth tolerance of +/- 3 seconds and a sequence length of 10.
@@ V Results and Discussion
This section presents and compares the performance of our event-based VPR system. Section V-A contains our evaluation of each ensemble strategy across reconstruction methods and feature extractors. Section V-B investigates the effect of our modified sequence matching approach (Section III-D) compared to the original SeqSLAM methodology. The influence of event-binning strategies on recognition accuracy and robustness is then examined in Section V-C. Finally, Section V-D compares key design choices in event-based VPR, including binning strategies, polarity, reconstruction methods, and feature extractors, critically evaluating the impact of each component on performance.
@@ V-A Event VPR via Ensembling
The effectiveness of the proposed ensemble strategies across reconstruction methods, feature extractors and the combination of feature extractors, reconstruction and temporal ensembles was evaluated against two baselines: (i) the best-performing individual ensemble member and (ii) the temporal resolution only ensemble from [11]. The results exhibit a bimodal distribution, reflecting the expected performance gap between day–day and day–night pairs due to significant domain shifts. To highlight these trends, results are grouped into day and night conditions. As shown in Figure 3, across 10 diverse traverse combinations, the proposed ensembles consistently outperformed both baselines.
Table I quantifies comparisons to existing methods along with the relative improvement over the temporal ensemble baseline [11]. Under severe illumination variation on BrisbaneEventVPR (day to night), the proposed ensembling improves average recall@1 (AR@1) from 1% for LENS [8] and 37.41% for Temporal Ensembles [11] to 44.16% and 66.39% for the ensemble of feature extractors and combined ensembles. On NSAVP day–night, the feature extractor and combined ensembles improve AR@1 from 28.40% to 32.18% and to 42.55%, respectively. Therefore, the combined ensembles have a relative improvement of up to 77% in day-night transitions. Gains in daytime conditions are smaller but consistent, with up to 3.02% relative gain on BrisbaneEventVPR and 8.63% relative gain on NSAVP. Additionally, figure 4 illustrates the fusion of predictions from multiple event-to-frame reconstructions, yielding a higher recall than the best individual method and highlights the robustness gained through aggregation.
TABLE I: Average Recall@1 and Gain over Temporal Ensemble [11].
Condition Method AR@1 (%) Gain (%)
D–D (Brisbane Event) LENS [8] 73.00 -11.23
Temporal Window [11] 82.24
Reconstruction 82.98 0.91
Feature Extractor 83.16 1.13
Combined Ensemble 84.72 3.02
D–N (Brisbane Event) LENS [8] 1.00 -97.33
Temporal Window [11] 37.41
Reconstruction 42.06 12.44
Feature Extractor 44.16 18.06
Combined Ensemble 66.39 77.48
D–D (NSAVP) Temporal Window [11] 56.88
Reconstruction 58.09 2.13
Feature Extractor 58.31 2.51
Combined Ensemble 61.79 8.63
D–N (NSAVP) Temporal Window [11] 28.40
Reconstruction 30.47 7.31
Feature Extractor 32.18 13.34
Combined Ensemble 42.55 49.83
Figure 3: Average Recall@1 (AR@1) for each reference–query pair across ensembling strategies and best individual method, shown separately for day and night conditions. Results are averaged over sequence lengths of 10, 20, and 30 at 1 Hz sampling. The combined ensemble aggregates predictions from varied feature extractors, event-to-frame reconstructions and temporal resolutions.
@@
TABLE II: AR@1 across sequence lengths and p-values from paired t-test
@@ V-C Event Binning Strategies
Figure 5: Average Recall@1 versus the number of reference and query frames, evaluated across two binning strategies and varying event slice resolutions. Time-based binning consistently yields the highest recall, with performance improving at higher resolutions (i.e., smaller bin sizes). The total number of frames serves as a proxy for computational cost, as both event-to-frame reconstruction and feature extraction scale linearly with frame count. Consequently, higher-resolution bins incur greater compute cost.
@@ V-D Reconstructions and Feature Extractors for Event VPR
Two variants of event count reconstruction are also evaluated: with and without polarity. Including polarity improves daytime performance by 6.8% Recall@1 (52.2% vs. 45.4%), while the difference at night is minimal (11.8% vs. 11.0%). Figure 8 shows VPR performance across different feature extractors. The observed trends are consistent with those reported in standard camera-based VPR benchmarks. Earlier methods, such as CosPlace and NetVLAD, achieve mean Recall@1 scores of 24–25%, while more recent approaches like MixVPR and MegaLoc reach 38–42%. MegaLoc, the current state-of-the-art, performs particularly well at night, achieving 16.9% Recall@1—at least 5% higher than all other evaluated methods.
V-E Computational Tradeoff
Table III reports the end-to-end runtime for a single query and AR@1 for different ensemble configurations, illustrating the accuracy versus compute tradeoff. The temporal-window [11] and reconstruction ensembles improve AR@1 by around 6–8% over the single model; while the feature extraction ensemble is most suitable for real-time application running at 13Hz with  10% increase in AR@1. The combined ensemble achieves the highest AR@1 but with substantially higher latency, making it most suitable for sparse global relocalisation or offline evaluation. A complementary correlation analysis of similarity matrices across reconstruction methods and feature extractors (included in our public code repository) shows that some representations, such as event count and event count without polarity, are strongly correlated, suggesting that future work could prioritise more complementary combinations to achieve better accuracy–runtime trade-offs.
TABLE III: Runtime vs Performance Analysis
Ensemble Type Runtime (ms) AR@1 (%)
None 47.63 47.07
Temporal Windows [11] 190.5 52.95
Event Reconstructions 190.5 54.61
VPR Feature Extractors 78.38 56.64
Combined Ensemble 1254.0 64.37
@@ VI Conclusions and Future Work
While our study includes a diverse set of feature extractors, reconstructions, and temporal strategies, further gains may be achieved by learning event-specific feature extractors and developing reconstructions tailored for VPR in night conditions. Notably, none of the deep learning-based feature extractors used were trained on event data, suggesting a promising direction for developing event-native architectures. Future work could also explore real-world deployment of our method, expanding the utility of event cameras in challenging environments.